OpenStudy (anonymous):

the length of a ectangle is 8m more than 6 tines the width the perietere is 156m find the width.. whaaaaat?? help?

OpenStudy (anonymous):

Can you rewrite your question without spelling mistakes please? My eyes are bleeding.

OpenStudy (anonymous):

sorryyy

OpenStudy (anonymous):

\[2x + 2y =156\]\[x =6y+8\]\[2(6y+8)+2y=156\]\[12y+2y +16 = 156\]\[14y = 140\]\[y=10\]\[2x + 20 = 156\]\[2x = 136\]\[x = 68\] Therefore the width and length are 10 m and 68m respectively